OpenStudy (anonymous):

find the partial derivatives fx (pi/3,1) if f(x,y)=x ln (y cos x)

OpenStudy (badhi):

consider y as a constant and differentiate wrt x, $$f_x(x,y)=\ln(y\cos x)+x\frac{-y\sin x}{y\cos x}=\ln(y\cos x)-x\tan x$$ now substitute x,y with pi/3 and 1 respectively $$f_x(\pi/3,1)=\ln(1.\cos(\pi/3))-\frac \pi3\tan\frac \pi 3=-\ln2-\frac \pi{\sqrt{3}}$$